# A Complex Scaling Method for Efficient and Accurate Scattering Emulation in Nuclear Reactions

### v1

@@ 3.3 No detection of anomaly
$V(r)=-V_{0}f(r;R,a)+\bm{\ell}\cdot\mathbf{s}\frac{V_{\mathrm{LS}}}{r}\frac{\mathrm{d}}{\mathrm{d}r}f(r;R,a),$ (14)
where $f$ is the Woods-Saxon function, and the parameter values are taken from Ref. [29]. This real potential is fitted to reproduce the $\mathrm{d}_{5/2}$ resonance at 1.274 MeV.
Previous studies reported that the KVP method exhibited an anomaly in the phase shifts of the $\mathrm{d}_{5/2}$ state at around 8.4 MeV in the center-of-mass frame. To investigate whether our method encounters any similar anomalies, we calculate and emulate the phase shift of this $\mathrm{d}_{5/2}$ state across very fine grids of center-of-mass energy. The parameter set we emulate is:
@@
We first select 15 sets of training parameters, $\{\bm{\Omega}_{i}\}_{i=1}^{15}$, with each value randomly generated within a ±$\pm 20\%$ interval of the best-fit values in Ref. [20]. Using these parameter sets, we apply principal component analysis to establish 8 reduced bases for the emulator.
To rigorously test the emulator’s accuracy, we perform the emulation calculation 20 times, each with different parameter sets, and subsequently compute the mean of the absolute relative errors defined in Eq. (13). The results are comprehensively displayed in Fig. 3, providing a clear visualization of the emulator’s performance across the energy spectrum.
@@
In the upper panel, the phase shifts are shown, with the black solid line representing the exact results and the red dashed line representing the emulation results. Notably, these two curves overlap completely and exhibit a smooth profile, indicating a high degree of reliability in our emulation method. The absence of abrupt changes or anomalies in these curves confirms the robustness of the complex scaling method used in our study. The lower panel shows the relative errors, which, despite their rapid fluctuations, primarily stabilize around $10^{-6}$. These fluctuations are attributed to the very fine energy grids used in our simulations. Nevertheless, the consistently low error magnitude underscores the emulator’s high precision. This level of accuracy, particularly with such small deviations from the exact results, demonstrates the effectiveness of our emulation approach in producing reliable and accurate phase shifts across a broad range of energies.

### v2

@@ 3.3 No detection of anomaly
$V(r)=-V_{0}\mathscr{Y}(r;R,a)+\bm{\ell}\cdot\mathbf{s}\frac{V_{\mathrm{LS}}}{r}\frac{\mathrm{d}}{\mathrm{d}r}\mathscr{Y}(r;R,a),$ (22)
where $\mathscr{Y}$ is the Woods-Saxon function, and the parameter values are taken from Ref. [29]. This real potential is fitted to reproduce the $\mathrm{d}_{5/2}$ resonance at 1.274 MeV.
Previous studies reported that the KVP method exhibited an anomaly in the phase shifts of the $\mathrm{d}_{5/2}$ state at around 8.4 MeV in the center-of-mass frame. To investigate whether our method encounters any similar anomalies, we calculate and emulate the phase shift of this $\mathrm{d}_{5/2}$ state across a center-of-mass energy range from 0.02 MeV to 10 MeV, using an energy step size of 0.02 MeV. The parameter set we emulate is:
@@
We select 15 sets of training parameters, $\{\bm{\Omega}_{i}\}_{i=1}^{15}$, with each value randomly generated within a ±$\pm 20\%$ interval of the best-fit values in Ref. [20], and we confine our parameter space within this range. Using exact solutions on these parameter sets, we apply principal component analysis to establish 8 reduced basis states for the emulator.
We initially emulate the phase shifts across the energy spectrum using the best-fit values from Ref. [20]. To rigorously evaluate the emulator’s accuracy, we perform the emulation calculations 20 times at each energy point, with each run employing a different parameter set randomly generated via Latin Hypercube Sampling (LHS) across the entire parameter space. We then compute the mean of the absolute relative errors at each energy point. The results are clearly visualized in Fig. 3, providing a comprehensive assessment of the emulator’s performance across the energy spectrum.
@@
In the upper panel, the phase shifts are shown, with the black solid line representing the exact results and the red dashed line representing the emulation results. Notably, these two curves overlap completely and exhibit a smooth profile, indicating a high degree of reliability in our emulation method. The absence of abrupt changes or anomalies in these curves confirms the robustness of the complex scaling method used in our study. The lower panel shows the mean of relative errors, which, despite their rapid fluctuations, primarily stabilize around $10^{-7}$. These fluctuations are attributed to the very fine energy grids used in our calculations. Nevertheless, the consistently low error magnitude underscores the emulator’s high precision. This level of accuracy, particularly with such small deviations from the exact results, demonstrates the effectiveness of our emulation approach in producing reliable and accurate phase shifts across a broad range of energies.